OpenStudy (angel_kitty12):

Can someone help me factor this equation? I just need you to explain some steps. Thanks. The equation is m^3-3m^2a+3ma^2-a^3

OpenStudy (angel_kitty12):

\[m^3-3m^2a+3ma^2-a^3\]

OpenStudy (anonymous):

factor by grouping

OpenStudy (angel_kitty12):

I have factored by grouping but i got confused on something. I'll show my work and explain what i'm confused on

OpenStudy (anonymous):

sounds good

OpenStudy (mathmale):

It just so happens that the expression you've shared is the cube of a binomial. Know where to find a formula for factoring that? Factoring by grouping should also work. In what way do you feel stuck?

OpenStudy (angel_kitty12):

\[(m^3-a^3) + (-3m^2a+3ma^2)\] \[(m-a) (m^2+ma+a^2)-3ma(m-a)\]

OpenStudy (angel_kitty12):

i have gotten stuck after this step

OpenStudy (angel_kitty12):

i believe that you square these two to get \[(m-a)^2 (m^2+ma+a^2)-3ma\]

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ (m-a) }} (m^2+ma+a^2)-3ma{\color{blue}{(m-a) }}\) notice any common factor?

OpenStudy (angel_kitty12):

this is where my confusion is. After this step, I'm unsure of what to do next.

OpenStudy (angel_kitty12):

can someone explain what i do after getting \[(m-a)^2 (m^2+ma+a^2) -3ma\]

OpenStudy (angel_kitty12):

I'm unsure of what to do with the -3ma

OpenStudy (jdoe0001):

well... hold the mayo

OpenStudy (jdoe0001):

but thus far, you're correct, one sec

OpenStudy (jdoe0001):

\({\color{blue}{ (m-a) }} (m^2+ma+a^2)-3ma{\color{blue}{(m-a) }} \\ \quad \\ (m-a)[(m^2+ma+a^2)-3ma] \\ \quad \\ (m-a)[m^2{\color{brown}{ -3ma+ma}}+a^2] \\ \quad \\ (m-a)[m^2-2ma+a^2] \\ \quad \\ \textit{now, recall what a perfect square trinomial is} \\ \quad \\ \begin{array}{cccccllllll} {\color{brown}{ a}}^2& + &2{\color{brown}{ a}}{\color{blue}{ b}}&+&{\color{blue}{ b}}^2\\ \downarrow && &&\downarrow \\ {\color{brown}{ a}}&& &&{\color{blue}{ b}}\\ &\to &({\color{brown}{ a}} + {\color{blue}{ b}})^2&\leftarrow \end{array}\qquad % perfect square trinomial, negative middle term \begin{array}{cccccllllll} {\color{brown}{ a}}^2& - &2{\color{brown}{ a}}{\color{blue}{ b}}&+&{\color{blue}{ b}}^2\\ \downarrow && &&\downarrow \\ {\color{brown}{ a}}&& &&{\color{blue}{ b}}\\ &\to &({\color{brown}{ a}} - {\color{blue}{ b}})^2&\leftarrow \end{array}\) and notice, that \(\bf (m-a)[{\color{purple}{ m^2-2ma+a^2 }} ]\impliedby {\color{purple}{ \textit{perfect square trinomial} }} \)

OpenStudy (angel_kitty12):

Question, where did the second (m-a) go in the equation though? I understand most of this but I don't understand why the (m-a) isn't squared.

OpenStudy (jdoe0001):

notice the 3rd line, the "brown" part, the addition

OpenStudy (jdoe0001):

m-a isn't squared? hmmm the first one? wasn't that the common factor you took from the second group? :)

OpenStudy (angel_kitty12):

No, that I understand. The -3ma +ma would equal to -2ma because of the 1 in front of the ma. But what I'm asking is why isn't the (m-a) and the (m-a) in -3ma together to make (m-a)^2

OpenStudy (angel_kitty12):

wait...now I'm a bit confused

OpenStudy (jdoe0001):

ohhhh why would they combine into the binomial at the 2nd power? well I gather you do not recall your "perfect square trinomial" then =) but usually comes out of FOIL, if you do say (m-a)(m-a), or \((m-a)^2\) you'll noitice you'll end up with \(m^2-2ma-a^2\)

OpenStudy (angel_kitty12):

wouldn't you get \[(m-a)^2 (m^2-2ma+a^2)\]

OpenStudy (angel_kitty12):

oh wait so the answer would be (m-a)^3

OpenStudy (angel_kitty12):

because it's a perfect square trinomial oh

OpenStudy (jdoe0001):

ahemm nope because the (m-a) on the left-side,was just the common factor of the second group and yes, \((m-a)(m-a)^2\implies (m-a)^3\)

OpenStudy (angel_kitty12):

but i usually thought that you had to break apart the trinomial \[(m^2-2ma+a^2) = (m-ma-a) (m+ma-a)\]

OpenStudy (angel_kitty12):

then you get (m-a) (m-a)

OpenStudy (angel_kitty12):

that's why i assumed it would be \[(m-a)^4\] because of the first (m-a)^2 and the two new (m-a)

OpenStudy (jdoe0001):

ohhh ahemmm not quite you're thinking \(\textit{difference of squares} \\ \quad \\ (a-b)(a+b) = a^2-b^2\qquad \qquad a^2-b^2 = (a-b)(a+b)\) that's not true for a perfect square trinomial \(\bf a^2-b^2 \ne (a-b)^2\)

OpenStudy (angel_kitty12):

wait, can you elaborate a bit more? I may have confused the two now. My teacher taught all the factoring at once and I may have mixed up the rules.

OpenStudy (angel_kitty12):

oh wait

OpenStudy (angel_kitty12):

so would that mean that (m-a)^2 would equal to (m-a)... no i think i confused myself

OpenStudy (jdoe0001):

well... hmm what part is throwing you off? \((m-a)^2 \implies (m-a)(m-a)\)

OpenStudy (angel_kitty12):

i just don't understand how the \[(m-a)^2\] ends up being (m-a) and why the answer is (m-a)^3 instead of (m-a)^4

OpenStudy (jdoe0001):

hmm.... lemme hmm hmmm do it from the start, shouldn't take long

OpenStudy (angel_kitty12):

Thank you so much

OpenStudy (jdoe0001):

\(\bf m^3-3m^2a+3ma^2-a^3 \\ \quad \\ (m^3-a^3)-(3m^2a-3ma^2) \\ \quad \\ (m-a)(m^2+ma+a^2)-(3m^2a-3ma^2) \\ \quad \\ {\color{blue}{ (m-a) }}(m^2+ma+a^2)-3ma{\color{blue}{ (m-a) }} \\ \quad \\ {\color{blue}{ (m-a)}}[(m^2+ma+a^2)-3ma] \\ \quad \\ {\color{blue}{ (m-a) }}[m^2{\color{brown}{ -3ma+ma}}+a^2] \\ \quad \\ {\color{blue}{ (m-a) }}[m^2-2ma+a^2] \\ \quad \\ {\color{blue}{ (m-a) }}[{\color{purple}{ m^2-2ma+a^2 }} ]\impliedby {\color{purple}{ \textit{perfect square trinomial} }} \\ \quad \\ {\color{blue}{ (m-a) }}(m-a)^2\implies (m-a)^3\)

OpenStudy (mathmale):

Once again: There is a basic formula for factoring the difference of two cubes, and you could save a lot of time by looking that up. Look up (e. g., through Google): special products and factors or "factoring the difference of two cubes" Alternatively, try assuming that (a-b) is a factor and divide that into a^3-3a^2*b-3ab^2-b^3.

OpenStudy (angel_kitty12):

Okay i think i understand more now. thank you all so much. I found my mistake.

OpenStudy (jdoe0001):

yw